OpenStudy (anonymous):

I'm crying in tears here... I just need help with this.. why won't anybody help me??? Find the number of units x that produces the minimum average cost per unit C(with line over the C) in the given equation. C = 0.001x^3 + 5x + 2

ganeshie8 (ganeshie8):

looks there is a typo in your equation, it should be : C = 0.001x^\(\color{red}{2}\) + 5x + 2 can you please check once ? :)

OpenStudy (anonymous):

No, it's x^3.. :(

OpenStudy (anonymous):

it would be easier if it was x^2... I just don't get this cost stuff and the c with a line over it.. it doesn't make sense to me

ganeshie8 (ganeshie8):

Oh yeah they want to minimize the average cost \(\large \overline{C}\)

OpenStudy (anonymous):

the example in the book has a weird derivative all over x.. i don't know why it does that for this type of problem.. why not just take the derivative and find critical points?

OpenStudy (anonymous):

yeah.. i didn't know how to do that symbol in this website but yeah, thats what it's looking for

ganeshie8 (ganeshie8):

can you please take a screenshot of example problem and attach ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh, hold on for example problem.. that was the problem im trying to solve.. :/

ganeshie8 (ganeshie8):

no its okay :) i have understood hte problem

ganeshie8 (ganeshie8):

lets try this : suppose you went for shopping and purchased 100 icecreams, and lets say you had spent $1000 for those icecreams can you tell me the cost of each icecream ?

OpenStudy (anonymous):

each one cost $10.00

ganeshie8 (ganeshie8):

yes, but how did u get $10.00 ?

OpenStudy (anonymous):

i divided

ganeshie8 (ganeshie8):

100 icecreams = $1000 so, 1 icecream = $1000/100

ganeshie8 (ganeshie8):

good :) if u had purchased "x" icecreams, and if it costs you "n" dollars, can you tell me the cost for each icecream ?

OpenStudy (anonymous):

umm, x/n?

ganeshie8 (ganeshie8):

think again

OpenStudy (anonymous):

n/x?

ganeshie8 (ganeshie8):

sure ?

OpenStudy (anonymous):

no...

ganeshie8 (ganeshie8):

n/x is correct, to be sure, visualize a scenario : x = 2 icecreams n = 10 dollars whats the cost of each icecream ?

ganeshie8 (ganeshie8):

is it 2/10 or 10/2 ?

OpenStudy (anonymous):

letters throw me off.. but i see it as 10/2

ganeshie8 (ganeshie8):

they throw everyone off haha we just need to get use to them by using them more :)

OpenStudy (anonymous):

right.. :)

ganeshie8 (ganeshie8):

okay so one more final question on icecreams to you : supppose you purchased "x" icecreams, and it costed you "0.001x^3 + 5x + 2" dollars whats the cost of each icecream ?

OpenStudy (anonymous):

(0.001x^3+5x+2)/x

ganeshie8 (ganeshie8):

what does that expression represent ?

OpenStudy (anonymous):

cost of icecream?

ganeshie8 (ganeshie8):

thats the cost per icecream, right ?

ganeshie8 (ganeshie8):

yes both are same

ganeshie8 (ganeshie8):

instead of icecreams, if we say "x" represents some other object/unit, then your expression represents "cost per unit" right ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

i hope you see why we're dividing "x" when we're asked to find `average cost per unit`

OpenStudy (anonymous):

so the c with line over it means average cost per unit..

ganeshie8 (ganeshie8):

Exactly ! \[\text{Average cost per item}(\overline{C}) = \dfrac{\text{Total cost of items}}{\text{number of items}}\]

OpenStudy (anonymous):

so how do i solve?..

ganeshie8 (ganeshie8):

divide the given expression by "x" then optimize it as usual

ganeshie8 (ganeshie8):

you know how to optimize a funciton, right ?

OpenStudy (anonymous):

umm... no.. :(

ganeshie8 (ganeshie8):

okay its easy if you know derivatives the method is same for optimizing ANY function

ganeshie8 (ganeshie8):

whats our function to be optimized ?

OpenStudy (anonymous):

so the whole thing would be C= .001x^3+5x+2/x So would C' be .001x^2+5+(2/x)?

ganeshie8 (ganeshie8):

exactly !

ganeshie8 (ganeshie8):

step1 : find the derivative, set it equal to 0, solve x

OpenStudy (anonymous):

i'm a little shaky on that.. and then to find x is harder for me.. :/

ganeshie8 (ganeshie8):

function to be optimized : \(\large f(x) = .001x^2+5+(2/x)\)

ganeshie8 (ganeshie8):

find the derivative : \(\large f'(x) = ?\)

OpenStudy (anonymous):

.... oh man.. these fractions.. how do i change fractions into something easier?

ganeshie8 (ganeshie8):

\(\large 2/x = 2x^{-1}\)

OpenStudy (anonymous):

would it be = 0.002x-2/x^2?

ganeshie8 (ganeshie8):

\[\large f(x) = .001x^2+5+(2/x)\] \[\large \begin{align} f'(x) &= .001(2x)+0+2(-1/x^2) \\ &= .002x - 2/x^2 \end{align} \]

ganeshie8 (ganeshie8):

Yep ! set it equal to 0 and solve x

OpenStudy (anonymous):

0.002x-2/x^2=0.. solving for x.. whats best way to do this? i always just plug in random numbers until i get close to 0 if i can.. but i don't know if thats really best way to go about it..

ganeshie8 (ganeshie8):

the kings way of solving it is to get rid off the fraction first : multiply x^2 both sides

ganeshie8 (ganeshie8):

\[\large \begin{align} f'(x) &= 0 \\ .002x - 2/x^2 &=0\\ 0.002x^3 - 2 &=0 \end{align} \]

ganeshie8 (ganeshie8):

I have just mulitplied x^2 both sides, is that clear so far ?

OpenStudy (anonymous):

yeah, but how come the .002 is x^3 now? do you have to multiply every term by x^2? because i would have just multiplied x^2 by x^2 and then x^2 by 0..

ganeshie8 (ganeshie8):

yes, instead of describing u how it works, let me show u more steps, then it will be clear : \[\large 0.002x - \dfrac{2}{x^2} = 0\] multiply x^2 both sides : \[\large x^2*\left(0.002x - \dfrac{2}{x^2}\right) = x^2*(0)\]

OpenStudy (anonymous):

gotcha :)

ganeshie8 (ganeshie8):

good :) so what do u get for x after u solve ?

OpenStudy (anonymous):

i divide both sides by .002 right?

ganeshie8 (ganeshie8):

actually i don't like decimals, so i would multiply by 1000 first

ganeshie8 (ganeshie8):

\[\large 0.002x^3-2 = 0\] multiply 1000 both sides, you get : \[\large 2x^3-2000 = 0\]

ganeshie8 (ganeshie8):

now divide 2 both sides, you get : \[\large x^3-1000 = 0\]

ganeshie8 (ganeshie8):

\[\large x^3 = 1000\] \[\large x = ?\]

OpenStudy (anonymous):

1000/3

ganeshie8 (ganeshie8):

nopes

OpenStudy (anonymous):

i don't know how to undo a cube...

OpenStudy (anonymous):

a square is square root right?.. but a cube..

ganeshie8 (ganeshie8):

a cube has got a cuberoot too

ganeshie8 (ganeshie8):

\[\large x^3 = 1000\] \[\large x^3 = 10^3\] \[\large x = ?\]

OpenStudy (anonymous):

soo... x=1000

ganeshie8 (ganeshie8):

nope x = 10

OpenStudy (anonymous):

i should be able to plug 10 in for x into C' right to get 0?

ganeshie8 (ganeshie8):

x=10 is the value that minimizes the average cost per unit

OpenStudy (anonymous):

holy cow! That is alot of algebra that i haven't used in, in forever... if at all.. I'm not a mathmatecian for sure!

OpenStudy (anonymous):

I really appreciate your patence with me and walking through all the tedious small steps that get me in tears.. i just was so upset tonight with trying to figure this out. I give you infinity medals!

ganeshie8 (ganeshie8):

np :) btw, you're good in math as you're asking lot of tough questions you won't be asking questions if you're not a mathematician haha!

ganeshie8 (ganeshie8):

have good sleep :)

OpenStudy (anonymous):

lol, i got 4 more questions before bed.. :) thanks for the help!!